OpenStudy (anonymous):

a cooler contains 7 cans of soda 3 colas, 3 oranges, and 1 cherry. Two cans are selected at random without replacement. Find the prob that at least one can is cola

OpenStudy (anonymous):

3/7*3/7=9/49

OpenStudy (anonymous):

scratch that

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

Wow i feel stupid 3/7*7/7 i think so 21/49 seems more reasonable.

OpenStudy (anonymous):

does that seem right?

OpenStudy (anonymous):

It makes sense the first time I did I was intending both cans to be cola.

OpenStudy (anonymous):

It seems more reasonable, but is still wrong.

OpenStudy (anonymous):

so how do u do it?

OpenStudy (anonymous):

Draw a probability tree!

OpenStudy (anonymous):

Well than Sir Newton provide your expertise.

OpenStudy (anonymous):

Well, your method give the answer 3/7, which is clearly wrong; that's as if only one was picked, and completely ignores what happens if you pick another one first THEN cola; try again, or draw a tree!

OpenStudy (anonymous):

3/7*4/7=12/49?

OpenStudy (anonymous):

The probability of picking one in two attempts isn't less than picking one in one attempt. Try again.

OpenStudy (anonymous):

I dont get it order doesnt matter

OpenStudy (anonymous):

You either pick one first (3/7) or you pick one of the others first (4/7). If you pick it first (3/7) you can ignore that because whatever you pick next will make it true. For the others (4/7), you now have a 3/6 chance of picking cola. It's (3/7) * 1 + (4/7) * (3/6)

OpenStudy (anonymous):

but how can u times those fractions?

OpenStudy (anonymous):

is it the 3/6?

OpenStudy (anonymous):

can u please help me

OpenStudy (anonymous):

(a/b)*(c/d) = (ac)/(bd)

OpenStudy (dumbcow):

15/21